Hector Company has developed the following standard costs for its product for 2019:

HECTOR COMPANY
Standard Cost Card
Product A
Cost Element Standard Quantity × Standard Price = Standard Cost
Direct materials 4 pounds $3 $12
Direct labor 3 hours 8 24
Manufacturing overhead 3 hours 4 12
$48


The company expected to produce 30,000 units of Product A in 2020 and work 90,000 direct labor hours.

Actual results for 2020 are as follows:
31,000 units of Product A were produced.
Actual direct labor costs were $746,200 for 91,000 direct labor hours worked.
Actual direct materials purchased and used during the year cost $346,500 for 126,000 pounds.
Actual variable overhead incurred was $155,000 and actual fixed overhead incurred was $205,000.


Compute the following variances showing all computations to support your answers. Indicate whether the variance are favorable or unfavorable?

(a) Materials Quantity Variance $

FavorableNot ApplicableUnfavorable

(b) Total Direct Labor Variance $

FavorableUnfavorableNot Applicable

(c) Direct Labor Quantity Variance $

Not ApplicableUnfavorableFavorable

(d) Direct Materials Price Variance $

FavorableUnfavorableNot Applicable

(e) Total Overhead Variance $

UnfavorableFavorableNot Applicable
Answer #1

a. Material quantity variance = (31000*4-126000)*3 = 6000 unfavorable

b. Total direct labor variance = 746200 - 31000*24 = 2200 favorable

c. Direct labor quantity variance = (31000*3-91000)*8 = 16000 favorable

d. Direct material price variance = (3-346500/126000)*126000 = 31500 favorable

e. Total overhead variance = 31000*12 - 155000-205000 = 12000 favorable